OpenStudy (anonymous):

How many lead cubes of sides 3cm could be made from a lead cube of side 27cm?

OpenStudy (anonymous):

volume of big lead cube divided by volume of small lead cube

OpenStudy (anonymous):

\[Answer = \frac{27^{3}}{3^{3}} = \frac{3^{9}}{3^{3}} = 3^{9 - 3} = 3^{6}\]

OpenStudy (anonymous):

Ok a lead cube would be a square right?

OpenStudy (anonymous):

yes. volume of a cube is side x side x side

OpenStudy (anonymous):

so that would be 27^3=19683 & 3^3= 27

OpenStudy (anonymous):

yes. now divide

OpenStudy (anonymous):

ok i got 729. Should this answer be written in cm^3?

OpenStudy (anonymous):

correct and yes

OpenStudy (anonymous):

awesome! thanks :)